# Pentagon relation in quantum cluster scattering diagrams

### v4

@@ 1. Introduction
In [GHKK18] the scattering diagram method was employed to study cluster algebras [FZ02], where all information of a cluster algebra or a cluster pattern is encoded in the corresponding cluster scattering diagram (CSD, for short). It was clarified implicitly in [GHKK18] and more explicitly in [Nak23] that the dilogarithm elements $\Psi[n]$, where $n$ is its normal vector, and the pentagon relation among them play prominent roles in a CSD.
@@
Meanwhile, it has been known that cluster algebras admit natural quantizations [BZ05, Tra11, FG09a, FG09b]. The relation to the quantum dilogarithm [FK94] was also studied in [FG09b, Kel11, KN11, Nag11]. Naturally, quantum cluster scattering diagrams (QCSD, for short), which are quantum analogues of CSDs, were introduced and studied in [Man21, DM21], where quantum dilogarithm elements again play a primary role. Here and below, we distinguish the quantum dilogarithm $\boldsymbol{\Psi}_{q}(x)$ and the quantum dilogarithm elements $\Psi_{a,b}[n]$ to avoid the confusion. The former is a formal power series of an indeterminate $x$ with a parameter $q$ due to [Sch53, FK94], while the latter are certain elements of the structure group $G$ of a QCSD. It is well known that the quantum dilogarithm satisfies the pentagon identity [FK94, CF99], which is a quantum analogue of the pentagon identity (more often called Abel’s identity or the five-term relation) of the dilogarithm [Lew81].
@@
Having this development in mind, in this paper we formulate the pentagon relation among the quantum dilogarithm elements (Theorem 2.7), which is the counterpart of the one for the dilogarithm elements for a CSD in [Nak23]. The definition of QCSDs is just as parallel to the one for CSDs thanks to a more general formulation of scattering diagrams by [KS14]. However, a quantum dilogarithm element $\Psi_{a,b}[n]$ carries some additional rational parameters (the quantum data) $a$ and $b$ compared with its classical counterpart $\Psi[n]$, and this is the source of the complication and the richness in the quantum case. (The situation is somewhat similar to the fusion of representations of quantum affine algebras with the spectral parameter.)
@@
As an application of the pentagon relation, together with the combinatorics of the quantum data, we establish the nonpositivity of a certain class of nonskew-symmetric QCSDs (Theorem 5.6 and Corollary 5.7), which extends the results of [LLRZ14, DM21, CFMM20]. Also, we explicitly present various consistency relations and their positivity for QCSDs of rank 2 completely (Section 3.2) or up to some degree (Section 5.4), most of which are new in the literature.
@@
Acknowledgment. We thank Peigen Cao for useful comments and especially for pointing out the error in the proof of Lemma 4.2 in the early manuscript. This work is partly supported by JSPS Grant Numbers JP16H03922 and JP22H01114.
@@ 2.1. Structure group $G$
Here we introduce the underlying structure group $G$ of a QCSD following [Man21, DM21, CFMM20, KS14]. See also [GHKK18, Nak23] for parallel notions for a CSD.
@@
So far, the only difference from the classical case is the Lie bracket in (2.5). It is a quantum analogue of the Lie bracket
@@ 2.2. $y$-representation
Let $y$ be a formal symbol. Let $\mathbb{Q}(q^{1/\delta_{0}})[y]_{q}$ be the noncommutative and associative algebra over $\mathbb{Q}(q^{1/\delta_{0}})$ with generators $y^{n}$ ($n\in N^{+}\sqcup\{0\})$ and the relations
@@
(2.8) $\sum_{n\in Q}c_{n}y^{n}\quad(c_{n}\in\mathbb{Q}(q^{1/\delta_{0}})).$
@@ 2.3. Quantum dilogarithm elements and pentagon relation
Let $\boldsymbol{\Psi}_{q}(x)$ be the quantum dilogarithm by [Sch53, FV93, FK94],
@@
The equality holds, because the both expressions satisfy the property
@@
where $n_{0}\in N_{\mathrm{pr}}^{+}$ is the one such that $n=jn_{0}$ for some $j\in\mathbb{Z}_{>0}$. We call it a quantum dilogarithm element. For simplicity, we also write
@@
We call the parameters $a$ and $b$ the quantum data of $\Psi_{a,b}[n]$. They control the interval and the shift of $\Psi_{a,b}[n]$, respectively, from the standard one $\Psi_{1,0}[n]$.
@@
The above $\Psi_{a,b}[n]$ is a quantum analogue of the dilogarithm element in the classical case [Nak23, §III.1.4]
@@
By (2.10), a dilogarithm element $\Psi_{a,b}[n]$ acts on $\mathcal{R}_{q}(y)$ under $\rho_{y}$ as
@@
The same formula in the RHS of (2.21) is also obtained by the adjoint action of the quantum dilogarithm $\boldsymbol{\Psi}_{q^{1/s\delta(n)}}(q^{b}y^{n})$ on $\mathcal{R}_{q}(y)$ [FG09a]. In this sense we have the correspondence between $\Psi_{1/s\delta(n),b}[n]$ and the quantum dilogarithm $\boldsymbol{\Psi}_{q^{1/s\delta(n)}}(q^{b}y^{n})$.
@@
(b). (Pentagon relation. Cf. [FK94, §2]). If $\{n_{2},n_{1}\}=c$ $(c\in(1/\delta_{0})\mathbb{Z}_{>0})$, we have, for any $b_{i}$,
@@
(2.26) $\begin{split}(\Psi_{c,b_{2}}[n_{2}]\Psi_{c,b_{1}}[n_{1}])(y^{n_{1}})&=\Psi_{c,b_{2}}[n_{2}](y^{n_{1}})\\ &=y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}),\end{split}$
@@
(2.27) $\begin{split}&\quad\ (\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}]\Psi_{c,b_{2}}[n_{2}])(y^{n_{1}})\\ &=(\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}])(y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}))\\ &=\Psi_{c,b_{1}}[n_{1}](y^{n_{1}}(1+q^{c+b_{1}+b_{2}}y^{n_{1}+n_{2}})\\ &\qquad\times(1+q^{c+b_{2}}y^{n_{2}}(1+q^{-c+b_{1}+b_{2}}y^{n_{1}+n_{2}})^{-1}))\\ &=\Psi_{c,b_{1}}[n_{1}](y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}+q^{c+b_{1}+b_{2}}y^{n_{1}+n_{2}}))\\ &=y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}(1+q^{-c+b_{1}}y_{1})^{-1}+q^{c+b_{1}+b_{2}}y^{n_{1}+n_{2}}(1+q^{-c+b_{1}}y_{1})^{-1})\\ &=y^{n_{1}}(1+q^{-c+b_{1}}y_{1}+q^{c+b_{2}}y^{n_{2}}+q^{b_{1}+b_{2}}y^{n_{2}}y^{n_{1}})(1+q^{-c+b_{1}}y_{1})^{-1}\\ &=y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}).\end{split}$
@@
(2.28) $\begin{split}&\quad\ (\Psi_{c,b_{2}}[n_{2}]\Psi_{c,b_{1}}[n_{1}])(y^{n_{2}})\\ &=\Psi_{c,b_{2}}[n_{2}](y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}})^{-1})\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}))^{-1}\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}}+q^{b_{1}+b_{2}}y^{n_{1}}y^{n_{2}})^{-1},\end{split}$
@@
(2.29) $\begin{split}&\quad\ (\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}]\Psi_{c,b_{2}}[n_{2}])(y^{n_{2}})\\ &=(\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}])(y^{n_{2}})\\ &=\Psi_{c,b_{1}}[n_{1}](y^{n_{2}}(1+q^{-c+b_{1}+b_{2}}y^{n_{1}+n_{2}})^{-1})\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}})^{-1}(1+q^{-c+b_{1}+b_{2}}y^{n_{1}+n_{2}}(1+q^{-c+b_{1}}y_{1})^{-1})^{-1}\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}}+q^{b_{1}+b_{2}}y^{n_{1}}y^{n_{2}})^{-1}.\end{split}$
@@
Two equalities (2.30) and (2.31) are equivalent, and they do not depend on the choice of the sign therein. We show (2.30) for the sign $+$ as follows:
(2.32) $\begin{split}\Psi_{a,b}[n]&=\exp\Biggl{(}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[ja]_{q}}\frac{q^{jpa}-q^{-jpa}}{q^{jpa}-q^{-jpa}}q^{bj}X_{jn}\Biggr{)}\\ &=\exp\Biggl{(}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[jpa]_{q}}\frac{q^{jpa}-q^{-jpa}}{q^{ja}-q^{-ja}}q^{bj}X_{jn}\Biggr{)}\\ &=\exp\Biggl{(}\sum_{t=1}^{p}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[jpa]_{q}}q^{j(2t-p-1)a}q^{bj}X_{jn}\Biggr{)}\\ &=\prod_{t=1}^{p}\Psi_{pa,b+(2t-p-1)a}[n].\end{split}$
@@ 3.1. Quantum cluster scattering diagrams
and $\gamma^{\prime}$ is the velocity vector of $\gamma$ at the wall $\mathbf{w}_{\lambda}$. The product $\mathfrak{p}_{\gamma,\mathfrak{D}}$ is an infinite one in general, and it is well-defined in $G$ due to the finiteness condition. We say that a pair of scattering diagrams $\mathfrak{D}$ and $\mathfrak{D}^{\prime}$ are equivalent if $\mathfrak{p}_{\gamma,\mathfrak{D}}=\mathfrak{p}_{\gamma,\mathfrak{D}^{\prime}}$ for any admissible curve $\gamma$ for both $\mathfrak{D}$ and $\mathfrak{D}^{\prime}$. We say that a scattering diagram $\mathfrak{D}$ is consistent if $\mathfrak{p}_{\gamma,\mathfrak{D}}=\mathrm{id}$ for any admissible loop (i.e., closed curve) $\gamma$ for $\mathfrak{D}$.
@@
To each admissible loop $\gamma$, the consistency condition $\mathfrak{p}_{\gamma,\mathfrak{D}}=\mathrm{id}$ gives a relation among quantum dilogarithm elements. Under the correspondence in Remark 2.6 this relation is translated as an identity for the quantum dilogarithm (quantum dilogarithm identity), which is a generalization of a quantum dilogarithm identity associated with a period of a quantum cluster pattern studied in [FG09b, Kel11, KN11, Nag11]. It is also the quantum counterpart of the (classical) dilogarithm identity for the Rogers dilogarithm in [Nak21].

### v7

@@ 1. Introduction
In [GHKK18] the scattering diagram method was employed to study cluster algebras [FZ02], where all information of a cluster algebra or a cluster pattern is encoded in the corresponding cluster scattering diagram (CSD, for short). It was clarified implicitly in [GHKK18] and more explicitly in [Nak23] that the dilogarithm elements $\Psi[n]$, where $n$ is its normal vector, and the pentagon relation among them play prominent roles in a CSD. Meanwhile, it has been known that cluster algebras admit natural quantizations [BZ05, Tra11, FG09a, FG09b], where the quantum dilogarithm plays a key role. Naturally, quantum cluster scattering diagrams (QCSD, for short), which are quantum analogs of CSDs, were introduced and studied in [Man21, DM21], where quantum dilogarithm elements play a primary role. Here and below, we distinguish the quantum dilogarithm $\boldsymbol{\Psi}_{q}(x)$ and the quantum dilogarithm elements $\Psi_{a,b}[n]$. The former is a formal power series of an indeterminate $x$ with a parameter $q$ due to [Sch53, FK94, FG09a], while the latter are certain elements of the structure group $G$ of a QCSD. It is well known that the quantum dilogarithm satisfies the pentagon identity [FK94, CF99], which is a quantum analog of the pentagon identity (more often called Abel’s identity or the five-term relation) of the dilogarithm [Lew81].
@@
Having this development in mind, in this paper we formulate the pentagon relation among the quantum dilogarithm elements (Theorem 2.7), which is the counterpart of the pentagon relation for the dilogarithm elements for a CSD in [Nak23]. The definition of QCSDs is in parallel with the one for CSDs thanks to a more general formulation of scattering diagrams by [KS14]. However, each quantum dilogarithm element $\Psi_{a,b}[n]$ carries some additional rational parameters (the quantum data) $a$ and $b$ compared with its classical counterpart $\Psi[n]$, and this is the source of the complication and the richness in the quantum case. (The situation is somewhat similar to the fusion of representations of quantum affine algebras with the spectral parameter.)
@@
As an application, we study the positivity of QCSDs in Section 5. We first formulate the positivity of CSDs and QCSDs (Definitions 5.1 and 5.4) based on the structure of the walls therein. It is known that any CSD is positive [GHKK18] and it implies the positivity of the theta functions for each CSD. It is also known that any QCSD with the skew-symmetric initial exchange matrix is positive [DM21]. On the other hand, there are examples of nonpositive theta functions for QCSDs [LLZ12, CFMM24]. This implies that QCSDs are not always positive in the nonskew-symmetric case. Using the pentagon relation, we clarify the origin of the nonpositivity in the combinatorics of quantum dilogarithm elements, and we establish the nonpositivity of a certain class of nonskew-symmetric QCSDs (Theorem 5.10 and Corollary 5.12). This extends the results of [LLRZ14, DM21, CFMM24] systematically. Also, we explicitly present various consistency relations and their positivity for QCSDs of rank 2 completely (Section 3.2) or up to degree 4 (Section 5.4). Most of these formulas are new in the literature and exhibit the complexity of the positivity phenomenon.
@@
Acknowledgment. We thank Peigen Cao for useful comments and especially for pointing out the errors in Section 4.3 in the early manuscript. This work is partly supported by JSPS Grant Numbers JP16H03922 and JP22H01114.
@@ 2.1. Structure group $G$
Here we introduce the underlying structure group $G$ of a QCSD following [KS14, Man21, DM21, CFMM24]. See also [GHKK18, Nak23] for parallel notions for a CSD.
@@
So far, the only difference from the classical case is the Lie bracket in (2.5). It is a quantum analog of the Lie bracket
@@ 2.2. $y$-representation
Let $y$ be a symbol. Let $\mathbb{Q}(q^{1/\delta_{0}})[y]_{q}$ be the noncommutative and associative algebra over $\mathbb{Q}(q^{1/\delta_{0}})$ with generators $y^{n}$ ($n\in N^{+}\sqcup\{0\})$ and the $q$-commutative elations
@@
(2.8) $\sum_{n\in N^{+}\sqcup\{0\}}c_{n}y^{n}\quad(c_{n}\in\mathbb{Q}(q^{1/\delta_{0}})).$
@@ 2.3. Quantum dilogarithm elements and pentagon relation
Let $\boldsymbol{\Psi}_{q}(x)$ be the quantum dilogarithm by [Sch53, FV93, FK94, FG09a],
@@
This convention is due to [FG09a]. The equality holds, because both expressions satisfy the property
@@
where $n_{0}\in N_{\mathrm{pr}}^{+}$ is the one such that $n=jn_{0}$ for some $j\in\mathbb{Z}_{>0}$. We call it a quantum dilogarithm element. We call the parameters $a$ and $b$ the quantum data of $\Psi_{a,b}[n]$. They control the interval and the shift in the sum, respectively. For simplicity, we also write
@@
The above $\Psi_{a,b}[n]$ is a quantum analog of the dilogarithm element in the classical case [Nak23, §III.1.4]
@@
By (2.10), each dilogarithm element $\Psi_{a,b}[n]$ acts on $\mathcal{R}_{q}(y)$ under $\rho_{y}$ as
@@
The same formula in the RHS of (2.21) is also obtained by the adjoint action of the quantum dilogarithm $\boldsymbol{\Psi}_{q^{1/s\delta(n)}}(q^{b}y^{n})$ on $\mathcal{R}_{q}(y)$ [FG09a]. In this sense, we have the correspondence between $\Psi_{1/s\delta(n),b}[n]$ and the quantum dilogarithm $\boldsymbol{\Psi}_{q^{1/s\delta(n)}}(q^{b}y^{n})$.
@@
(b). (Pentagon relation. Cf. [FK94, §2]). Suppose that $\{n_{2},n_{1}\}=c$ $(c\in(1/\delta_{0})\mathbb{Z}_{>0})$. Then, for any $b_{1}$ and $b_{2}$, we have
@@
$\begin{split}(\Psi_{c,b_{2}}[n_{2}]\Psi_{c,b_{1}}[n_{1}])(y^{n_{1}})&=\Psi_{c,b_{2}}[n_{2}](y^{n_{1}})\\ &=y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}),\end{split}$
@@
$\begin{split}&\quad\ (\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}]\Psi_{c,b_{2}}[n_{2}])(y^{n_{1}})\\ &=(\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}])(y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}))\\ &=\Psi_{c,b_{1}}[n_{1}](y^{n_{1}}(1+q^{c+b_{1}+b_{2}}y^{n_{1}+n_{2}})\\ &\qquad\times(1+q^{c+b_{2}}y^{n_{2}}(1+q^{-c+b_{1}+b_{2}}y^{n_{1}+n_{2}})^{-1}))\\ &=\Psi_{c,b_{1}}[n_{1}](y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}+q^{c+b_{1}+b_{2}}y^{n_{1}+n_{2}}))\\ &=y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}(1+q^{-c+b_{1}}y_{1})^{-1}+q^{c+b_{1}+b_{2}}y^{n_{1}+n_{2}}(1+q^{-c+b_{1}}y_{1})^{-1})\\ &=y^{n_{1}}(1+q^{-c+b_{1}}y_{1}+q^{c+b_{2}}y^{n_{2}}+q^{b_{1}+b_{2}}y^{n_{2}}y^{n_{1}})(1+q^{-c+b_{1}}y_{1})^{-1}\\ &=y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}).\end{split}$
@@
$\begin{split}&\quad\ (\Psi_{c,b_{2}}[n_{2}]\Psi_{c,b_{1}}[n_{1}])(y^{n_{2}})\\ &=\Psi_{c,b_{2}}[n_{2}](y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}})^{-1})\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}}(1+q^{c+b_{2}}y^{n_{2}}))^{-1}\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}}+q^{b_{1}+b_{2}}y^{n_{1}}y^{n_{2}})^{-1},\end{split}$
@@
$\begin{split}&\quad\ (\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}]\Psi_{c,b_{2}}[n_{2}])(y^{n_{2}})\\ &=(\Psi_{c,b_{1}}[n_{1}]\Psi_{c,b_{1}+b_{2}}[n_{1}+n_{2}])(y^{n_{2}})\\ &=\Psi_{c,b_{1}}[n_{1}](y^{n_{2}}(1+q^{-c+b_{1}+b_{2}}y^{n_{1}+n_{2}})^{-1})\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}})^{-1}(1+q^{-c+b_{1}+b_{2}}y^{n_{1}+n_{2}}(1+q^{-c+b_{1}}y_{1})^{-1})^{-1}\\ &=y^{n_{2}}(1+q^{-c+b_{1}}y^{n_{1}}+q^{b_{1}+b_{2}}y^{n_{1}}y^{n_{2}})^{-1}.\end{split}$
@@
Two equalities (2.26) and (2.27) are equivalent. The equality (2.26) is shown as follows:
$\begin{split}\Psi_{a,b}[n]&=\exp\Biggl{(}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[ja]_{q}}\frac{q^{jpa}-q^{-jpa}}{q^{jpa}-q^{-jpa}}q^{bj}X_{jn}\Biggr{)}\\ &=\exp\Biggl{(}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[jpa]_{q}}\frac{q^{jpa}-q^{-jpa}}{q^{ja}-q^{-ja}}q^{bj}X_{jn}\Biggr{)}\\ &=\exp\Biggl{(}\sum_{t=1}^{p}\sum_{j=1}^{\infty}\frac{(-1)^{j+1}}{j[jpa]_{q}}q^{j(2t-p-1)a}q^{bj}X_{jn}\Biggr{)}\\ &=\prod_{t=1}^{p}\Psi_{pa,b+(2t-p-1)a}[n].\end{split}$
@@ 3.1. Quantum cluster scattering diagrams
and $\gamma^{\prime}$ is the velocity vector of $\gamma$ at the wall $\mathbf{w}_{\lambda}$. The product $\mathfrak{p}_{\gamma,\mathfrak{D}}$ is infinite in general, and it is well-defined in $G$ due to the finiteness condition. We say that a pair of scattering diagrams $\mathfrak{D}$ and $\mathfrak{D}^{\prime}$ are equivalent if $\mathfrak{p}_{\gamma,\mathfrak{D}}=\mathfrak{p}_{\gamma,\mathfrak{D}^{\prime}}$ for any admissible curve $\gamma$ for both $\mathfrak{D}$ and $\mathfrak{D}^{\prime}$. We say that a scattering diagram $\mathfrak{D}$ is consistent if $\mathfrak{p}_{\gamma,\mathfrak{D}}=\mathrm{id}$ for any admissible loop (i.e., closed curve) $\gamma$ for $\mathfrak{D}$.
@@
To each admissible loop $\gamma$, the consistency condition $\mathfrak{p}_{\gamma,\mathfrak{D}}=\mathrm{id}$ gives a relation among quantum dilogarithm elements. Under the correspondence in Remark 2.6 this relation is translated as an identity for the quantum dilogarithm (quantum dilogarithm identity), which is a generalization of a quantum dilogarithm identity associated with a period of a quantum cluster pattern studied in [FG09b, Kel11, KN11, Nag11]. It is also the quantum counterpart of the (classical) dilogarithm identity for the Rogers dilogarithm in [Nak24].